Find the distance between the points to the nearest tenth. L(-4,11),M(-3,4)
1
Expert's answer
2011-10-11T10:35:30-0400
Let's calculate the distance between these points.

D = ((x1-x2)^2 + (y1-y2)^2)^0.5 = ((-4-(-3))^2 + (11-4)^2)^0.5 = (1 + 49)^0.5 = 50^0.5 = 7,07106...
Rounding to the nearest tenth, D = 7,0.